# Sometimes Two Irrational Guards are Needed

### v3

@@ 1 Introduction
The art gallery problem was formulated in 1973 by Victor Klee. See, for example, the book by O’Rourke [46, page 2]. One of the earliest results states that every simple polygon on $n$ vertices can always be guarded with $\lfloor n/3\rfloor$ guards [22, 31].
@@
Figure 2: Any triangulation of a simple polygon can be three-colored. At least one of the color classes has at most $\lfloor n/3\rfloor$ vertices. This color class also guards the entire polygon, as every triangle is incident to all three colors [31].
@@
Interestingly, it is actually very tough to find any positive algorithmic results on the art gallery problem. It seems like the art gallery problem is almost impenetrable. For instance, only in 2002, Micha Sharir pointed out that the problem was even decidable [26, 27, see acknowledgments]. The decidability of the art gallery problem is actually easy once you know methods from real algebraic geometry [5]. The idea is to reduce the problem to the first-order theory of the reals. We encode guard positions by variables, and then we check if every point in the polygon is seen by at least one guard. Note that this is easy to encode in the first-order theory of the reals, as we are allowed to use existential ($\exists g_{1},g_{2},\ldots$) and universal quantifiers ($\forall p=(x,y)$). Since then, despite much research on the art gallery problem, no better algorithm appeared, as far as worst-case complexity is concerned. The underlying reason for the difficulty to find better algorithms can be explained by the fact that the art gallery problem is $\exists\mathbb{R}$-complete [57, 3]. In a nutshell, $\exists\mathbb{R}$-completeness precisely entails that there is no better method for the worst-case complexity of the problem. ($\exists\mathbb{R}$ can be defined as the class of problems that are equivalent to finding a real root to a multivariate polynomial with integer coordinates. See Section˜1.3 for an introduction.) More specifically, it was shown that arbitrary algebraic numbers may be needed to describe an optimal solution to the art gallery problem. This may come as a surprise to some readers, and was clearly a surprise back then. Specifically, “in practice”, it seems very rare that irrational guards are ever needed. The reason is that a typical situation is one of the following two. Either the guards have some freedom to move around and still see the entire polygon. Or if a guard has no freedom, it is forced to be on a line defined by vertices of the polygon. As the vertices of the polygon are at rational coordinates, the guards will be at rational coordinates in that case as well. Indeed, only in 2017, the first polygon requiring irrational guards was found [2]. Even though $\exists\mathbb{R}$-reductions exhibit an infinite number of polygons that require irrational guards, those polygons are not “concrete” in the naive sense of the word. And up to this day, this is the only “concrete” polygon [2] that we know does require irrational guards. In this work, we find a second polygon. It is superior to the first one in the sense that it shows that two guards are already enough to enforce irrational guards. As a single guard can always be chosen to have rational coordinates, we settle the question of the minimum number of guards required to have irrational guards.
@@ 1.1 Discussion
It is known that one guard can always be chosen to be rational [40]. The polygon by Abrahamsen, Adamaszek, and Miltzow [2] requires three irrational guards. The main strength of our finding is to determine the minimum number of guards required to have irrational guards.
@@
One way to circumvent worst-case complexity is to discretize the polygon and restrict oneself to a dense grid [15, 26, 27]. The polygon by Abrahamsen, Adamaszek, and Miltzow showed that a grid cannot have a better approximation factor than $4/3=1.333\ldots$. We improve this lower bound to $3/2=1.5$.
@@
Note that Bonnet and Miltzow showed that under some mild assumptions, the grid contains a constant factor approximation [15].
@@
The principal methods that we used for the construction of our polygon are in spirit similar to the methods used by Abrahamsen, Adamaszek, and Miltzow [2]. However, it turned out that it was considerably more difficult to find the polygon. On the one hand, our construction is smaller and thus there were fewer parameters that we had to manipulate to find a solution. On the other hand, the two guards interact in more intertwined ways. Thus making it much harder to find a correct placement of all the polygon vertices.
To be concrete, the construction by [2] has the guards $a$, $m$, and $t$. Guards $a$ and $m$ cover together two pockets. Similarly, guards $m$ and $t$ cover together two separate pockets. Thus there is no direct interaction between guards $a$ and $t$. This makes it easier to construct the different parts independently.
@@
Both our polygon and the polygon by [2] have their guards in the interior. It is an interesting open problem if one can enforce irrational guards, in case all guards are restricted to lying on the boundary and are only required to guard the boundary.
@@
One may wonder whether it is possible to enforce irrational guards on polygons with some extra properties like being rectilinear or monotone. Both of those questions have been positively resolved by [2].
@@ 1.2 Art Gallery Problem
In 1979, the first algorithm for guarding a polygon in linear time with one guard appeared [40]. It took until 1992 until there was an algorithm that could determine if a polygon could be guarded by two guards in $O(n^{4})$ time [7, 6]. As mentioned already above it took until 2002 to find the first correct algorithm that solves the art gallery problem [26, 27]. There is still no other algorithm known.
On the lower bound side, we know NP-hardness [39], APX-hardness [28] and W[1]-hardness [16]. One may argue that NP-hardness is enough evidence that there are no efficient algorithms for the art gallery problem and that this may fully explain the lack of algorithmic results. However, for other NP-complete problems like Clique, Subset-sum, Dominating Set, and TSP, we do know a myriad of algorithms. Although many of them run in exponential time in the worst case they give huge improvements in many different situations. We believe that the lack of algorithmic results may stem from the fact that we do not know how to discretize the art gallery problem efficiently. Note that all the mentioned problems are already discrete. We believe that the $\exists\mathbb{R}$-completeness of the art gallery problem may give the most compelling explanation of why a concise discretization of the art gallery problem is unlikely [3, 57]. Specifically, many discretization schemes would imply that the art gallery problem lies in NP and thus imply $\text{NP}=\exists\mathbb{R}$. The first proof that the art gallery problem is $\exists\mathbb{R}$-comlete was given by Abrahamsen, Adamaszek, and Miltzow in 2017 [3]. It was recently improved by Stade, who showed that it is even $\exists\mathbb{R}$-complete if we only require the boundary to be guarded [57]. It remains open whether guarding the boundary from the boundary is $\exists\mathbb{R}$-complete.
@@
There is a series of papers that studied the art gallery problem from a practical perspective. In other words, they implemented algorithms and tested them on benchmark instances [17, 32, 35, 59, 34]. The practical experiences from those papers suggest that irrational coordinates do not play any role in the pursuit to find an optimal solution.
@@
We are aware of two theoretical explanations for the discrepancy between the theoretical and the practical results. One such finding uses smoothed analysis and argues that there is with high probability an optimal solution after a small random perturbation of the polygon [34, 30]. A second explanation comes from Hengeveld and Miltzow. They introduced the notion of vision-stability. To explain this concept, we consider guards that can either see by some small angle $\delta$ around reflex vertices or are blocked by an angle $\delta$ by reflex vertices. See the green visibility regions in Figure˜4.
@@
This question was positively answered by Bertschinger, El Maalouly, Miltzow, Schnider, and Weber for homotopy-equivalence [10] and shortly improved by Stade and Tucker-Foltz to homeomorphic-equivalence [58].
@@ 1.3 Existential theory of the Reals
Here the first inclusion follows because a SAT instance can trivially be written as an equivalent ETR instance. The second inclusion is highly non-trivial and was first proven by Canny in his seminal paper [19].
Note that the complexity of working with continuous numbers was studied in various contexts. To avoid confusion, let us make some remarks on the underlying machine model. The underlying machine model for $\exists\mathbb{R}$ (over which sentences need to be decided and where reductions are performed) is the word RAM (or equivalently, a Turing machine) and not the real RAM [30] or the Blum-Shub-Smale model [14].
The complexity class $\exists\mathbb{R}$ gains its importance by numerous important algorithmic problems that have been shown to be complete for this class in recent years. The name $\exists\mathbb{R}$ was introduced by Schaefer in [48] who also pointed out that several NP-hardness reductions from the literature actually implied $\exists\mathbb{R}$-hardness. For this reason, several important $\exists\mathbb{R}$-completeness results were obtained before the need for a dedicated complexity class became apparent.
Common features of $\exists\mathbb{R}$-complete problems are their continuous solution space and the nonlinear relations between their variables. Important $\exists\mathbb{R}$-completeness results include the realizability of abstract order types [45, 56] and geometric linkages [49], as well as the recognition of geometric segment graphs [38, 42], unit-disk graphs [37, 43], and ray intersection graphs [20]. More results appeared in the graph drawing community [25, 29, 41, 50], regarding polytopes [24, 47], the study of Nash-equilibria [8, 11, 12, 33, 51], training neural networks [4, 9], matrix factorization [21, 53, 54, 55], or continuous constraint satisfaction problems [44]. In computational geometry, we would like to mention the art gallery problem [3, 57] and covering polygons with convex polygons [1].
Recently, the community started to pay more attention to higher levels of the “real polynomial hierarchy”, which surprisingly captures some interesting algorithmic problems [13, 23, 25, 36, 52, 18].
@@ 2.2 Guard Segments
We can force a guard to be positioned on a line segment within the polygon. Such a line segment is called a guard segment. Guard segments are commonly used in the context of the art gallery problem [2, 57]. In this section, we will describe how we construct a guard segment. Let us denote by $s$ the segment and by $\ell$ its supporting line.
@@ 2.3 Guarding Quadrilateral Pockets
Figure 8: Our complete polygon. The art gallery is shaded according to the function of each region: gray is the core, yellow is the pockets used to create guard segments, and turquoise are other pockets. The yellow dashed lines represent the guard segments. The coordinates of important vertices are given.
@@ 4 Challenges
We encountered new challenges while searching for our polygon, compared to Abrahamsen, Adamaszek, and Miltzow [2]’s polygon that requires three irrational guards.
@@
Allowing for universal quantifiers, we get the following simple formula [27]:
@@
The art gallery problem has a much simpler formulation when we allow universal quantifiers as seen above. So, another approach to find an ETR-formula of reasonable length for our instance of the art gallery problem is to use the formulation with universal quantifiers and perform quantifier elimination using some known algorithm. However, despite the formulation only using four existential quantifiers and two universal quantifiers we aim to eliminate, the resulting formula is seemingly infeasibly large as we will see. To illustrate this, we make some “back-of-the-envelope” calculation. There are several algorithms readily available in the literature; we use the algorithm from Basu et al. [5] as this is one of the most modern treatments of algorithms in real algebraic geometry. The results could vary for other algorithms. To give an upper bound on the running time, we denote with $k$ the numbers of existential quantifiers, with $l$ the number of universal quantifiers that we want to eliminate, with $s$ the number of polynomials, and $d$ denotes the degree of the polynomials. Clearly, $k=4$, and $l=2$. Furthermore, as the formula deals with lines and determinants only, $d=2$. The upper bound is given by
@@
In comparison to the polygon construction with three guards [2], our polygon has fewer parameters that we can adjust, as we have one guard less. As everything depends on those few parameters, it was difficult to find a configuration that satisfies all the desired properties simultaneously. Furthermore, we need to check some additional properties that didn’t play a role in the previous construction, as the middle guard was surrounded by the other guards from two different sides. At last, we could not avoid the supporting line of the guard segment intersecting two quadrilateral pockets.
Acknowledgments.
We would like to thank Thekla Hamm and Ivan Bliznets for their helpful comments on the presentation. Lucas Meijer is generously supported by the Netherlands Organisation for Scientific Research (NWO) under project no. VI.Vidi.213.150. Tillmann Miltzow is generously supported by the Netherlands Organisation for Scientific Research (NWO) under project no. 016.Veni.192.250 and no. VI.Vidi.213.150.

### v4

@@ Abstract
keywords:
Art Gallery Problem, Existential Theory of the Reals, Irrational Numbers
@@ 1 Introduction
The art gallery problem was formulated in 1973 by Victor Klee. See, for example, the book by O’Rourke (O’Rourke, 1987, page 2). One of the earliest results states that every simple polygon on $n$ vertices can always be guarded with $\lfloor n/3\rfloor$ guards Chvátal (1975); Fisk (1978).
@@
Figure 2: Any triangulation of a simple polygon can be three-colored. At least one of the color classes has at most $\lfloor n/3\rfloor$ vertices. This color class also guards the entire polygon, as every triangle is incident to all three colors Fisk (1978).
@@
Interestingly, it is actually very tough to find any positive algorithmic results on the art gallery problem. It seems like the art gallery problem is almost impenetrable. For instance, only in 2002, Micha Sharir pointed out that the problem was even decidable (Efrat and Har-Peled, 2002, 2006, see acknowledgments). The decidability of the art gallery problem is actually easy once you know methods from real algebraic geometry Basu et al. (2006). The idea is to reduce the problem to the first-order theory of the reals. We encode guard positions by variables, and then we check if every point in the polygon is seen by at least one guard. Note that this is easy to encode in the first-order theory of the reals, as we are allowed to use existential ($\exists g_{1},g_{2},\ldots$) and universal quantifiers ($\forall p=(x,y)$). Since then, despite much research on the art gallery problem, no better algorithm appeared, as far as worst-case complexity is concerned. The underlying reason for the difficulty to find better algorithms can be explained by the fact that the art gallery problem is $\exists\mathbb{R}$-complete Stade (2025); Abrahamsen et al. (2022). In a nutshell, $\exists\mathbb{R}$-completeness precisely entails that there is no better method for the worst-case complexity of the problem. ($\exists\mathbb{R}$ can be defined as the class of problems that are equivalent to finding a real root to a multivariate polynomial with integer coordinates. See Section˜1.3 for an introduction.) More specifically, it was shown that arbitrary algebraic numbers may be needed to describe an optimal solution to the art gallery problem. This may come as a surprise to some readers, and was clearly a surprise back then. Specifically, “in practice”, it seems very rare that irrational guards are ever needed. The reason is that a typical situation is one of the following two. Either the guards have some freedom to move around and still see the entire polygon. Or if a guard has no freedom, it is forced to be on a line defined by vertices of the polygon. As the vertices of the polygon are at rational coordinates, the guards will be at rational coordinates in that case as well. Indeed, only in 2017, the first polygon requiring irrational guards was found Abrahamsen et al. (2017). Even though $\exists\mathbb{R}$-reductions exhibit an infinite number of polygons that require irrational guards, those polygons are not “concrete” in the naive sense of the word. And up to this day, this is the only “concrete” polygon Abrahamsen et al. (2017) that we know does require irrational guards. In this work, we find a second polygon. It is superior to the first one in the sense that it shows that two guards are already enough to enforce irrational guards. As a single guard can always be chosen to have rational coordinates, we settle the question of the minimum number of guards required to have irrational guards.
@@ 1.1 Discussion
It is known that one guard can always be chosen to be rational Lee and Preparata (1979). The polygon by Abrahamsen, Adamaszek, and Miltzow Abrahamsen et al. (2017) requires three irrational guards. The main strength of our finding is to determine the minimum number of guards required to have irrational guards.
@@
One way to circumvent worst-case complexity is to discretize the polygon and restrict oneself to a dense grid Bonnet and Miltzow (2017); Efrat and Har-Peled (2002, 2006). The polygon by Abrahamsen, Adamaszek, and Miltzow showed that a grid cannot have a better approximation factor than $4/3=1.333\ldots$. We improve this lower bound to $3/2=1.5$.
@@
Note that Bonnet and Miltzow showed that under some mild assumptions, the grid contains a constant factor approximation Bonnet and Miltzow (2017).
@@
The principal methods that we used for the construction of our polygon are in spirit similar to the methods used by Abrahamsen, Adamaszek, and Miltzow Abrahamsen et al. (2017). However, it turned out that it was considerably more difficult to find the polygon. On the one hand, our construction is smaller and thus there were fewer parameters that we had to manipulate to find a solution. On the other hand, the two guards interact in more intertwined ways. Thus making it much harder to find a correct placement of all the polygon vertices.
To be concrete, the construction by Abrahamsen et al. (2017) has the guards $a$, $m$, and $t$. Guards $a$ and $m$ cover together two pockets. Similarly, guards $m$ and $t$ cover together two separate pockets. Thus there is no direct interaction between guards $a$ and $t$. This makes it easier to construct the different parts independently.
@@
Both our polygon and the polygon by Abrahamsen et al. (2017) have their guards in the interior. It is an interesting open problem if one can enforce irrational guards, in case all guards are restricted to lying on the boundary and are only required to guard the boundary.
@@
One may wonder whether it is possible to enforce irrational guards on polygons with some extra properties like being rectilinear or monotone. Both of those questions have been positively resolved by Abrahamsen et al. (2017).
@@ 1.2 Art Gallery Problem
In 1979, the first algorithm for guarding a polygon in linear time with one guard appeared Lee and Preparata (1979). It took until 1992 until there was an algorithm that could determine if a polygon could be guarded by two guards in $O(n^{4})$ time Belleville (1992, 1991). As mentioned already above it took until 2002 to find the first correct algorithm that solves the art gallery problem Efrat and Har-Peled (2002, 2006). There is still no other algorithm known.
On the lower bound side, we know NP-hardness Lee and Lin (1986), APX-hardness Eidenbenz et al. (2001) and W[1]-hardness Bonnet and Miltzow (2020). One may argue that NP-hardness is enough evidence that there are no efficient algorithms for the art gallery problem and that this may fully explain the lack of algorithmic results. However, for other NP-complete problems like Clique, Subset-sum, Dominating Set, and TSP, we do know a myriad of algorithms. Although many of them run in exponential time in the worst case they give huge improvements in many different situations. We believe that the lack of algorithmic results may stem from the fact that we do not know how to discretize the art gallery problem efficiently. Note that all the mentioned problems are already discrete. We believe that the $\exists\mathbb{R}$-completeness of the art gallery problem may give the most compelling explanation of why a concise discretization of the art gallery problem is unlikely Abrahamsen et al. (2022); Stade (2025). Specifically, many discretization schemes would imply that the art gallery problem lies in NP and thus imply $\text{NP}=\exists\mathbb{R}$. The first proof that the art gallery problem is $\exists\mathbb{R}$-comlete was given by Abrahamsen, Adamaszek, and Miltzow in 2017 Abrahamsen et al. (2022). It was recently improved by Stade, who showed that it is even $\exists\mathbb{R}$-complete if we only require the boundary to be guarded Stade (2025). It remains open whether guarding the boundary from the boundary is $\exists\mathbb{R}$-complete.
@@
There is a series of papers that studied the art gallery problem from a practical perspective. In other words, they implemented algorithms and tested them on benchmark instances Borrmann et al. (2013); Friedrichs (2012); Hengeveld and Miltzow (2021b); Tozoni et al. (2016); Hengeveld and Miltzow (2021a). The practical experiences from those papers suggest that irrational coordinates do not play any role in the pursuit to find an optimal solution.
@@
We are aware of two theoretical explanations for the discrepancy between the theoretical and the practical results. One such finding uses smoothed analysis and argues that there is with high probability an optimal solution after a small random perturbation of the polygon Hengeveld and Miltzow (2021a); Erickson et al. (2020). A second explanation comes from Hengeveld and Miltzow. They introduced the notion of vision-stability. To explain this concept, we consider guards that can either see by some small angle $\delta$ around reflex vertices or are blocked by an angle $\delta$ by reflex vertices. See the green visibility regions in Figure˜4.
@@
This question was positively answered by Bertschinger, El Maalouly, Miltzow, Schnider, and Weber for homotopy-equivalence Bertschinger et al. (2022) and shortly improved by Stade and Tucker-Foltz to homeomorphic-equivalence Stade and Tucker-Foltz (2023).
@@ 1.3 Existential theory of the Reals
Here the first inclusion follows because a SAT instance can trivially be written as an equivalent ETR instance. The second inclusion is highly non-trivial and was first proven by Canny in his seminal paper Canny (1988). Interestingly, there are also oracle separation results indicating that those three complexity classes are distinct Hamm et al. (2025).
Note that the complexity of working with continuous numbers was studied in various contexts. To avoid confusion, let us make some remarks on the underlying machine model. The underlying machine model for $\exists\mathbb{R}$ (over which sentences need to be decided and where reductions are performed) is the word RAM (or equivalently, a Turing machine) and not the real RAM Erickson et al. (2020) or the Blum-Shub-Smale model Blum et al. (1989).
The complexity class $\exists\mathbb{R}$ gains its importance by numerous important algorithmic problems that have been shown to be complete for this class in recent years. The name $\exists\mathbb{R}$ was introduced by Schaefer in Schaefer (2010) who also pointed out that several NP-hardness reductions from the literature actually implied $\exists\mathbb{R}$-hardness. For this reason, several important $\exists\mathbb{R}$-completeness results were obtained before the need for a dedicated complexity class became apparent.
Common features of $\exists\mathbb{R}$-complete problems are their continuous solution space and the nonlinear relations between their variables. Important $\exists\mathbb{R}$-completeness results include the realizability of abstract order types Mnëv (1988); Shor (1991) and geometric linkages Schaefer (2013), as well as the recognition of geometric segment graphs Kratochvíl and Matoušek (1994); Matousek (2014), unit-disk graphs Kang and Müller (2012); McDiarmid and Müller (2013), and ray intersection graphs Cardinal et al. (2018). More results appeared in the graph drawing community Dobbins et al. (2022); Erickson (2019); Lubiw et al. (2022); Schaefer (2021), regarding polytopes Dobbins et al. (2019); Richter-Gebert and Ziegler (1995), the study of Nash-equilibria Berthelsen and Hansen (2019); Bilò and Mavronicolas (2016, 2017); Garg et al. (2018); Schaefer and Štefankovič (2017), training neural networks Abrahamsen et al. (2021); Bertschinger et al. (2023), matrix factorization Chistikov et al. (2016); Schaefer and Štefankovič (2018); Shitov (2018, 2017), or continuous constraint satisfaction problems Miltzow and Schmiermann (2022). In computational geometry, we would like to mention the art gallery problem Abrahamsen et al. (2022); Stade (2025) and covering polygons with convex polygons Abrahamsen (2022). A complete overview can be found in a recent compendium Schaefer et al. (2024).
Recently, the community started to pay more attention to higher levels of the “real polynomial hierarchy”, which surprisingly captures some interesting algorithmic problems Blanc and Hansen (2021); D’Costa et al. (2021); Dobbins et al. (2022); Jungeblut et al. (2022); Schaefer and Štefankovič (2023); Bürgisser and Cucker (2009). And even to the first order theory of the reals as a complexity class.
@@ 2.2 Guard Segments
We can force a guard to be positioned on a line segment within the polygon. Such a line segment is called a guard segment. Guard segments are commonly used in the context of the art gallery problem Abrahamsen et al. (2017); Stade (2025). In this section, we will describe how we construct a guard segment. Let us denote by $s$ the segment and by $\ell$ its supporting line.
@@ 2.3 Guarding Quadrilateral Pockets
Figure 8: Our complete polygon. Note that $v_{6}$ is cropped. The art gallery is shaded according to the function of each region: gray is the core, yellow is the pockets used to create guard segments, and turquoise are other pockets. The yellow dashed lines represent the guard segments. The coordinates of important vertices are given.
@@ 4 Challenges
We encountered new challenges while searching for our polygon, compared to Abrahamsen, Adamaszek, and Miltzow Abrahamsen et al. (2017)’s polygon that requires three irrational guards.
@@
Allowing for universal quantifiers, we get the following simple formula Efrat and Har-Peled (2006):
@@
The art gallery problem has a much simpler formulation when we allow universal quantifiers as seen above. So, another approach to find an ETR-formula of reasonable length for our instance of the art gallery problem is to use the formulation with universal quantifiers and perform quantifier elimination using some known algorithm. However, despite the formulation only using four existential quantifiers and two universal quantifiers we aim to eliminate, the resulting formula is seemingly infeasibly large as we will see. To illustrate this, we make some “back-of-the-envelope” calculation. There are several algorithms readily available in the literature; we use the algorithm from Basu et al. Basu et al. (2006) as this is one of the most modern treatments of algorithms in real algebraic geometry. The results could vary for other algorithms. To give an upper bound on the running time, we denote with $k$ the numbers of existential quantifiers, with $l$ the number of universal quantifiers that we want to eliminate, with $s$ the number of polynomials, and $d$ denotes the degree of the polynomials. Clearly, $k=4$, and $l=2$. Furthermore, as the formula deals with lines and determinants only, $d=2$. The upper bound is given by
@@
In comparison to the polygon construction with three guards Abrahamsen et al. (2017), our polygon has fewer parameters that we can adjust, as we have one guard less. As everything depends on those few parameters, it was difficult to find a configuration that satisfies all the desired properties simultaneously. Furthermore, we need to check some additional properties that did not play a role in the previous construction, as the middle guard was surrounded by the other guards from two different sides. At last, we could not avoid the supporting line of the guard segment intersecting two quadrilateral pockets.
Acknowledgements.
We would like to thank Thekla Hamm and Ivan Bliznets for their helpful comments on the presentation.